# Audio-Visual Target Speaker Extraction with Reverse Selective Auditory Attention

### v1

@@ IV-A2 VoxCeleb2
For in-domain evaluation, our system is trained on VoxCeleb2 or LRS2 and then evaluated on the corresponding test set.
IV-B Evaluation Set
@@ IV-D Baseline methods
IV-D1 AV-DPRNN
AV-DPRNN is the audio-visual dual-path RNN-based TSE system. Compared to SEANet in Figure 3, the difference is that there is no suppressor, pre-suppressor and speech-noise interactor in the AV-DPRNN. As shown in Figure 6, it predicts the target speech only so the performance gap can reflect the effect of reverse selective auditory attention. The loss function considers the output of all $(R+1)$ extractors. That setting is equivalent to the architecture of SEANet. AV-DPRNN has 4.12M parameters.
@@ IV-D3 AV-Sepformer
AV-Sepformer [6] (AV-Sep) is the transformer-based multimodal extraction framework motivated by the transformer-based audio-only speech separation system. It introduces the audio-visual interaction to learn the relationship between the audio and visual modality. AV-Sepformer also contains a dual-path structure to understand inter-chunk and intra-chunk knowledge. Also, it was only studied on VoxCeleb2 for the original paper, and we extended the experiments for the LRS2 dataset. Considering that AV-Sepformer has a very large model architecture, for relatively fair comparison and GPU memory consideration, our study slightly reduced the number of layers in AV-Sepformer. However, it still contains 21.73M parameters, which is around twice than that in SEANet.
IV-E Metrics
@@
Following the previous works, we use scale-invariant signal-to-distortion ratio (SI-SDR[48] and signal-to-distortion ratio (SDR) as the main metrics to evaluate the quality of the extracted speech. Their unit is dB. Meanwhile, we introduce several additional metrics to comprehensively judge the quality of the extraction: Perceptual evaluation of speech quality (PESQ[53] and short-term objective intelligibility (STOI[54] are introduced to precept the extracted speech. Furthermore, we also report these metrics with the relative improvement (denoted as the suffix ‘$i$’, i.e., SI-SDRi, SDRi, PESQi and STOIi). For instance, denote ground-truth speech as $s$, mixed audio as $x$, output speech as $\hat{s}$, there is $SDRi=SDR(\hat{s},s)-SDR(x,s)$. Higher is better for these 8 metrics. Moreover, we also compute the word-error-rate (WER) by feeding the extracted speech into the pre-trained automatic speech recognition systems [55]. It can represent the readability and the informativeness of the extracted speech. Lower is better for WER.
V Results and analysis
@@
This section reports the results with five subsections: Section V-A compares SEANet with the baseline methods and other previous methods for in-domain evaluation. Both the results on LRS2 and VoxCeleb2 dataset have been provided. Section V-B reports the cross-domain evaluation result for SEANet and the baseline methods. The evaluation sets are LRS3, Grid and TCD-TIMIT datasets. In Section V-C, we conduct the ablation study to verify the efficiency of the architecture of SEANet, especially for reverse selective auditory attention. Section V-D employs the quality study to prove the robustness of SEANet with different evaluation conditions and settings. Finally, section V-E provides some visualization results.
V-A In-domain evaluation: LRS2 and VoxCeleb2
@@
For in-domain evaluation, we conduct the experiments of SEANet and re-implement three pre-mentioned baselines with the same training setting for comparison. For LRS2 dataset training and evaluation, we report all nine metrics for four methods, for the experiments on the VoxCeleb2 dataset, WER cannot be reported since there is no public text labels. Meanwhile, we also provide the results of other previous methods based on the results in their published paper and compare them with SEANet.
V-A1 Comparison with baseline methods on the LRS2
@@
TABLE I: Comparison of SEANet and baseline methods for nine evaluation metrics, we train and evaluate them on the LRS2 dataset with the same setting. Note that all results come from our own re-implement for fair comparison. Lower is better for WER and higher is better for the rest metrics.
@@
V-A2 Comparison with baseline methods on the VoxCeleb2
Table II reports the performance of SEANet and baseline methods on the VoxCeleb2 dataset. Our SEANet performs the best SI-SDR and SDR on the VoxCeleb2, for instance, 12.88 dB SI-SDR from SEANet vs 10.84 SI-SDR from AV-Sep. Also, SEANet achieves the best PESQ and STOI with 2.34 and 0.89, respectively. Noted that the PESQ in VoxCeleb2 is slightly better than that in LRS2 because the videos in VoxCeleb2 have the longer duration. In summary, SEANet achieves an obvious improvement over the previous methods for all eight metrics. Considering that VoxCeleb2 contains the many noisy videos, these results verify that SEANet can handle the AV-TSE task well for the challenging dataset by analysing the noise intelligently.
TABLE II: Comparison of SEANet and baseline methods for eight evaluation metrics, we train and evaluate them on the VoxCeleb2 dataset with the same setting. Note that all results come from our own re-implement for fair comparison. Higher is better for all these metrics.
@@ V-A2 Comparison with baseline methods on the VoxCeleb2
V-A3 Comparison with other methods on the LRS2 or VoxCeleb2
@@
V-B Cross-domain evaluation: LRS3, Grid and TCD-TIMIT
@@
V-C Ablation study for efficiency
This section conducts the ablation studies to prove the efficiency of SEANet architecture and analyse why SEANet can achieve better performance. Here, we consider AV-DPRNN for comparison since it shares the similar model architecture with our SEANet, the difference is that SEANet involves reverse attention to learn the speech-noise interaction. As we mentioned, some videos in VoxCeleb2 are out of sync by -0.08 to 0.08 seconds. Our ablation studies are conducted on the LRS2 dataset.
V-C1 Reverse attention
@@
To verify that the proposed reverse selective auditory attention ability is important to assist the extraction process, we consider the following structures in Figure 7 for comparison: AV-DPRNN has the speech branch only. S1 contains the speech-speech interaction with attention (Eq. 2 contains self-attention $A_{s+}$ only). S2 has the noisy branch, the speech-noise interaction is conducted based on traditional self-attention (Eq. 2 contains $A_{s-}$ only without reverse operation). S3 is similar to S2, while speech-noise interaction based on reverse attention (Eq. 2 contains $A_{s-}$ only). S4 has speech-noise and speech-speech interaction attention, both of them apply traditional self-attention ($A_{s-}$ in Eq. 2 has no reverse operation). SEANet contains the reverse attention for speech-noise interaction and self-attention for speech-speech interaction.
@@
TABLE V: Ablation study for the different network structures. $\oplus$ denote typical self-attention, which is positive; $\ominus$ represents our proposed reverse attention, which is negative. Experiments are conducted on the LRS2 dataset. Illustrations of these structures can be found in Fig 7
@@ V-C2 Model size
AV-DPRNN 4.12M 10.24 10.81 $R=6,D=64$
-deeper 9.32M 11.25 11.80 $R=15,D=64$
-wider 11.88M 11.45 12.01 $R=6,D=128$
SEANet 8.70M 13.08 13.67 $R=6,D=64$
V-C3 Extraction correctness
@@
V-D Quality study for robustness
In this section, we conduct quality studies to learn the robustness of the proposed SEANet under the different training or evaluation settings. Similarly, we evaluate the AV-DPRNN system for comparison on the LRS2 dataset.
V-D1 Hyper-parameter
@@
V-D2 AV-TSE for ordinary acoustic environment
Then, we report the results of SEANet for ordinary acoustic environment: when the mixture of audio contains the speech from the target speaker and the non-speech noise only, our task can be viewed as audio-visual speech enhancement. We train the AV-DPRNN and SEANet based on these speech enhancement format data for comparison. In Table VII, we report the results with the different Signal-to-Noise Ratio (SNR) for the input audio. SEANet performs better than AV-DPRNN for this enhancement task under different noise levels with the very stable improvement. Meanwhile, a very high SI-SDR can be achieved in this enhancement task. Because extracting the target speaker’s speech from the non-speech noise is easier than extracting it from the mixture of speech.
@@ V-D2 AV-TSE for ordinary acoustic environment
TABLE VII: Performance of AV-DPRNN and the proposed SEANet in the ordinary acoustic environment for audio-visual speech enhancement task. The input audio contains the target speaker’s speech and the non-speech noise. Input SNR (dB) indicates the SNR for the input audio to represent the different noisy level.
V-D3 AV-TSE for challenging acoustic environments
Our previous results have shown that SEANet can handle the scenario of two speakers talking simultaneously. Here, we further study the robustness of our SEANet in challenging acoustic environments. As shown in Table VIII, the audio mixture contains the target speaker’s speech. Then, ‘S’ denotes the existence of one interference speaker. ‘N’ means the existence of non-speech noise, such as background noise or music. For instance, ‘S + S + N’ represents the input audio containing the target speaker’s voice, two other speakers’ voices and the non-speech noise. This can simulate very complex extraction conditions.
@@ V-D3 AV-TSE for challenging acoustic environments
TABLE VIII: Performance of AV-DPRNN and the proposed SEANet in the challenging acoustic environment. For the mixture of audio which contains the target speaker’s speech, ‘S’ denotes the existence of one interference speaker, and ‘N’ denotes the existence of the non-speech noisy signal.
V-E Visualization
@@ V-E1 Extraction visualization
Figure 11: Visualization for the attention score matrix $A_{s}$ in each PSNL block, the left and right column represents the intra-chunk and inter-chunk attention, respectively. The attention map for intra-chunk correlation focuses on the diagonals. In contrast, the attention map for inter-chunk correlation concentrates on the rows and columns.
V-E2 Attention score visualization
Figure 11 indicates the attention visualization for $A_{s}$ in Eq. 2. There are ten attention maps since SEANet contains $R=5$ PSNL blocks with intra-chunk and inter-chunk attention. Intra-chunk attention is used in each chunk to search the local information, it focuses on the diagonals frame-wise correlation. Inter-chunk attention works across all chunks to learn the global correlation, it focuses on the knowledge within rows and columns. The deeper layer learns more representative information. Each attention score contains the self-attention part to study the voice consistency and the reverse attention to exclude the noisy signals.
VI Discussion
We further discuss the problems we found in the experimental process and provide our analysis.
VI-A Extraction vs separation
@@ VI-B Video quality for AV-TSE
VII Conclusion
In this paper, we propose the Subtraction-and-ExtrAction network (SEANet) for audio-visual target speaker extraction (AV-TSE). SEANet focuses on reverse selective auditory attention to force the neural network to understand the noisy signal during extraction. Our experimental results verify that this ‘subtraction’ strategy can suppress noisy signal and avoid incorrect speaker extraction. SEANet outperforms previous methods for five different datasets under up to nine metrics. In future work, we will extend the reverse selective auditory attention mechanism to the audio-only target speaker extraction. Furthermore, we will analyse the roles of each auxiliary reference to achieve an intelligent balance between them.

### v3

@@ V-A2 VoxCeleb2
For in-domain evaluation, our system is trained on VoxCeleb2 or LRS2 and then evaluated on the corresponding test set. Note that in all our experiments, the audio mixtures are simulated based on two speech signals, except for the experiment described in Table X.
V-B Evaluation Set
@@ V-D Baseline methods
V-D1 AV-DPRNN
AV-DPRNN is the audio-visual dual-path RNN-based TSE system. Compared to SEANet in Figure 3, the difference is that there is no suppressor, pre-suppressor and speech-noise interactor in the AV-DPRNN. As shown in Figure 7, it predicts the target speech only so the performance gap can reflect the effect of reverse attention. The loss function considers the output of all $(R+1)$ extractors. That setting is equivalent to the architecture of SEANet. AV-DPRNN has 4.12M parameters.
@@ V-D3 AV-Sepformer
AV-Sepformer [6] (AV-Sep) is the transformer-based multimodal extraction framework motivated by the transformer-based audio-only speech separation system. It introduces the audio-visual interaction to learn the relationship between the audio and visual modality. AV-Sepformer also contains a dual-path structure to understand inter-chunk and intra-chunk knowledge. Also, it was only studied on VoxCeleb2 for the original paper, and we extended the experiments for the LRS2 dataset. Considering that AV-Sepformer has a very large model architecture, for relatively fair comparison and GPU memory consideration, our study slightly reduced the number of Transformer layers in AV-Sepformer from 8 to 6. However, it still contains 21.73M parameters, which is around twice than that in SEANet.
V-E Metrics
@@
Following the previous works, we use scale-invariant signal-to-distortion ratio (SI-SDR[50] and signal-to-distortion ratio (SDR) as the main metrics to evaluate the quality of the extracted speech. Their unit is dB. Meanwhile, we introduce several additional metrics to comprehensively judge the quality of the extraction: Perceptual evaluation of speech quality (PESQ[55] and short-term objective intelligibility (STOI[56] are introduced to precept the extracted speech. Furthermore, we also report these metrics with the relative improvement (denoted as the suffix ‘$i$’, i.e., SI-SDRi, SDRi, PESQi and STOIi). For instance, denote ground-truth speech as $s$, mixed audio as $x$, output speech as $\hat{s}$, there is $SDRi=SDR(\hat{s},s)-SDR(x,s)$. Higher is better for these 8 metrics. Moreover, we also compute the word-error-rate (WER) by feeding the extracted speech into the Whisper model (large version), which is the automatic speech recognition model [57] that pre-trained on 680k hours of diverse audio datasets. It can represent the readability and the informativeness of the extracted speech. Lower is better for WER.
VI Results and analysis
@@
This section reports the results with five subsections: Section VI-A compares SEANet with the baseline methods and other previous methods for in-domain evaluation. Both the results on LRS2 and VoxCeleb2 dataset have been provided. Section VI-B reports the cross-domain evaluation result for SEANet and the baseline methods. The evaluation sets are LRS3, Grid and TCD-TIMIT datasets. In Section VI-C, we conduct the ablation study to verify the efficiency of the architecture of SEANet, especially for reverse attention. Section VI-D employs the quality study to prove the robustness of SEANet with different evaluation conditions and settings. Finally, section VI-F provides some visualization results.
VI-A In-domain evaluation: LRS2 and VoxCeleb2
@@
In Table I and II, we conduct the experiments of SEANet and re-implement three pre-mentioned baselines with the same training setting for strictly fair comparison. For Table III, we provide the performance of other previous methods based on the results in their published paper and compare them with SEANet. Table I, II and III are used together to achieve a comprehensive and fair comparison with all previous solutions.
VI-A1 Comparison with baseline methods on the LRS2
@@
TABLE I: Comparison of SEANet and baseline methods for nine evaluation metrics, we train and evaluate them on the LRS2 dataset with the same setting. All results come from our own re-implement for fair comparison. Lower is better for WER and higher is better for the rest metrics.
@@
VI-A2 Comparison with baseline methods on the VoxCeleb2
Table II reports the performance of SEANet and baseline methods on the VoxCeleb2 dataset. Our SEANet performs the best SI-SDR and SDR on the VoxCeleb2, for instance, 12.88 dB SI-SDR from SEANet vs 10.84 SI-SDR from AV-Sep. Also, SEANet achieves the best PESQ and STOI with 2.34 and 0.89, respectively. Noted that the PESQ in VoxCeleb2 is slightly better than that in LRS2 because the videos in VoxCeleb2 have the longer duration. In summary, SEANet achieves an obvious improvement over the previous methods for all eight metrics.
TABLE II: Comparison of SEANet and baseline methods for eight evaluation metrics, we train and evaluate them on the VoxCeleb2 dataset with the same setting. All results come from our own re-implement for fair comparison. Higher is better for all these metrics.
@@ VI-A2 Comparison with baseline methods on the VoxCeleb2
VI-A3 Comparison with other methods on the LRS2 or VoxCeleb2
@@
VI-B Cross-domain evaluation: LRS3, Grid and TCD-TIMIT
@@
VI-C Ablation study for efficiency
This section conducts ablation studies to prove the efficiency of the proposed architecture and analyse why it can achieve better performance. We performed experiments on various architectures that were modified based on our framework or re-implemented from previous studies for comparison. Here, we consider AV-DPRNN for comparison since it shares the similar model architecture with our SEANet, the difference is that SEANet involves reverse attention to learn the speech-noise interaction. As we mentioned, some videos in VoxCeleb2 are out of sync by -0.08 to 0.08 seconds. Our ablation studies are conducted on the LRS2 dataset, since it is a high-quality audio-visual dataset with clean target speech.
VI-C1 Reverse attention
@@
To verify that the proposed reverse attention is important to assist the extraction process, we consider the following structures in Figure 8 for comparison: AV-DPRNN has the speech branch only. S1 contains the speech-speech interaction with attention (Eq. 2 contains self-attention $A_{s+}$ only). S2 has the noisy branch, the speech-noise interaction is conducted based on traditional self-attention (Eq. 2 contains $A_{s-}$ only without reverse operation). S3 is similar to S2, while speech-noise interaction based on reverse attention (Eq. 2 contains $A_{s-}$ only). S4 has speech-noise and speech-speech interaction attention, both of them apply traditional self-attention ($A_{s-}$ in Eq. 2 has no reverse operation). SEANet contains the reverse attention for speech-noise interaction and self-attention for speech-speech interaction.
@@
We further involve the previous method as baseline for comparison. [31, 30] also involve the noise during speech enhancement with a straightforward CNN-based interaction solution. We re-implement their approach in our AV-TSE system to achieve a fair comparison. As shown in Table V, SEANet performs better than [31, 30] since it explores more potential of speech-noise interaction.
TABLE V: Ablation study for the different network structures. $\oplus$ denote typical self-attention, which is positive; $\ominus$ represents our proposed reverse attention, which is negative. Experiments are conducted on the LRS2 dataset. We also re-implement the previous method in AV-TSE as the baseline for comparision. Illustrations of these structures can be found in Figure 8
@@ VI-C2 Model size
AV-DPRNN 4.12M 10.24 10.81 $R=5,D=64$
-deeper 9.32M 11.25 11.80 $R=14,D=64$
-wider 11.88M 11.45 12.01 $R=5,D=128$
SEANet 8.70M 13.08 13.67 $R=5,D=64$
VI-C3 Extraction correctness
@@
VI-C4 Model architecture
We further conducted experiments for SEANet with three different model architectures. As shown in Table VII, SEANetα performs much worse than SEANet due to the lack of extractor and suppressor (13.08 dB VS 10.57 dB). With pre-extractor and pre-suppressor only , SEANet can hardly learn high-quality speech representation. Meanwhile, SEANetβ performs slightly worse than SEANet, which means the noise-noise self-attention is useful in SEANet to guide the system to learn the noisy representation (13.08 dB VS 12.69 dB). That can further boost the speech representation through the speech-noise interaction block. SEANetγ modifies the order of subtraction and softmax in Eq 2:
$A_{s}=softmax{(\frac{(Q_{s}-Q^{\prime}_{n})\cdot{K_{s}}^{T}}{\sqrt{D}}})$ (8)
It achieves the similar result with SEANet (13.08 dB VS 12.98 dB). That verifies both implementation approaches are useful here. These experiments prove the reasonableness of SEANet.
TABLE VII: Ablation study for the PSNL block on the LRS2 dataset. Compared with SEANet, SEANetα denotes to remove the extractor and suppressor in PSNL block, SEANetβ denotes to remove the noise-noise self-attention in the suppressor, SEANetγ denotes to do subtraction on the logits first then feed into softmax.
Method SI-SDR SDR Remark
SEANet 13.08 13.67 -
SEANetα 10.57 11.17 Remove extractor/suppressor
SEANetβ 12.69 13.31 Remove noise-noise self-attention
SEANetγ 12.98 13.54 Do subtraction before softmax
VI-C5 Model efficiency
We further study the complexity of the SEANet model and compare it with two popular structures: CNN-based MuSE [5] and Transformer-based AV-Sepformer [6]. As shown in Table VIII. Our SEANet achieves the best performance with the relative small number of parameters and lower FLOPs. Furthermore, we conduct the experiments to study the relationship among model size, complexity and the performance in SEANet. SEANets and SEANetl denote the small version SEANet and large version SEANet, respectively.
TABLE VIII: Ablation study for the model complexity on the LRS2 dataset, $R$ and $D$ denote the depth of layers and dimension of layers in the SEANet, respectively.
Method # Param FLOPs SI-SDR Setting
MuSE [5] 15.01M 34.23G 10.97 -
AV-Sepformer [6] 21.73M 73.09G 12.60 -
SEANetD- 3.32M 8.46G 10.71 $R=5,D=32$
SEANetR- 5.12M 9.75G 11.25 $R=2,D=64$
SEANet 8.70M 21.17G 13.08 $R=5,D=64$
SEANetR+ 12.28M 36.59G 13.69 $R=8,D=64$
SEANetD+ 30.10M 72.72G 14.68 $R=5,D=128$
VI-D Quality study for robustness
In this section, we conduct quality studies to learn the robustness of the proposed SEANet under the different training or evaluation settings. Similarly, we evaluate the AV-DPRNN system for comparison on the LRS2 dataset since it contains the clean target speech.
VI-D1 Hyper-parameter
@@
VI-D2 AV-TSE for ordinary acoustic environment
Then, we report the results of SEANet for ordinary acoustic environment: when the mixture of audio contains the speech from the target speaker and the non-speech noise only, our task can be viewed as audio-visual speech enhancement. We train the AV-DPRNN and SEANet based on these speech enhancement format data for comparison. The non-speech noisy signals are sampled from the MUSAN dataset [62], which contains the nature noises (the sounds from thunder, rain, ball, etc) and the background music (playing the instrument or singing). The SNR is randomly selected from -5 dB to 5 dB. In Table IX, we report the results with the different Signal-to-Noise Ratio (SNR) for the input audio. SEANet performs better than AV-DPRNN for this enhancement task under different noise levels with the very stable improvement. Meanwhile, a very high SI-SDR can be achieved in this enhancement task. Because extracting the target speaker’s speech from the non-speech noise is easier than extracting it from the mixture of speech.
TABLE IX: Performance of AV-DPRNN and the proposed SEANet in the ordinary acoustic environment for audio-visual speech enhancement task. The input audio contains the target speaker’s speech and the non-speech noise. Input SNR (dB) indicates the SNR for the input audio to represent the different noisy level.
@@ VI-D2 AV-TSE for ordinary acoustic environment
VI-D3 AV-TSE for diverse acoustic environments
Our previous results have shown that SEANet can handle the scenario of two speakers talking simultaneously. Here, we further study the robustness of our SEANet in diverse acoustic environments on the LRS2 dataset. As shown in Table X, the audio mixture contains the target speaker’s speech. Then, ‘S’ denotes the existence of one interference speaker. ‘N’ means the existence of non-speech noise, such as background noise or music. For instance, ‘S + S + N’ represents the input audio containing the target speaker’s voice, two other speakers’ voices and the non-speech noise. This can simulate very complex extraction conditions.
@@ VI-D3 AV-TSE for diverse acoustic environments
TABLE X: Performance of AV-DPRNN and the proposed SEANet in the challenging acoustic environment. For the mixture of audio which contains the target speaker’s speech, ‘S’ denotes the existence of one interference speaker, and ‘N’ denotes the existence of the non-speech noisy signal.
@@
VI-E Multi-modal interaction
In this section, we provide the results of the proposed multi-modal interaction method on the LRS2 dataset. All experiments are fine-tuned on the pre-trained SEANet structure to speed up the training process.
As shown in Table XI, all three methods get robust performance for SI-SDR and SDR. The best one is F-SEANet, which achieves 13.63 dB SI-SDR, P-SEANet and A-SEANet achieve 13.36 dB SI-SDR and 13.51 dB SI-SDR, respectively. These results show that involving vision information in the early stage can benefit the system more, while the late stage of involvement can hardly boost the system.
We propose the new assumptions to the importance of visual information in AV-TSE. In the long-term research process, one consensus is that lip movements can guide the extraction system in understanding the detailed target speech representation along the time axis. However, the lip movements may play the role of the ‘mentor’ to hint the system knows which kind of voice should be extracted with a global view. The evidence is that it is more important to involve lip movements in the early stage of the extraction network than in the late stage. In the future, we will focus more on the efficient design in the early-stage audio-visual fusion to confirm whether our arguments are correct.
TABLE XI: Ablation study for the audio-visual interaction learning based on SEANet. The results are conducted on the LRS2 dataset.
Method SI-SDR SDR Remark
AV-DPRNN 10.24 10.81 -
MuSE 10.97 11.57 -
AV-Sep 12.60 13.13 -
SEANet 13.08 13.67 -
F-SEANet 13.63 14.20 Fusion-based
P-SEANet 13.36 13.93 PSNL-based
A-SEANet 13.51 14.08 Algorithm-based
VI-F Visualization
@@
VII Discussion
VII-A Extraction vs separation
@@ VII-B Video quality for AV-TSE
VIII Conclusion
In this paper, we propose the Subtraction-and-ExtrAction network (SEANet) for audio-visual target speaker extraction (AV-TSE). SEANet focuses on selective auditory attention to force the neural network to understand the noisy signal during extraction. Our experimental results verify that this ‘subtraction’ strategy can suppress noisy signal and avoid incorrect speaker extraction. SEANet outperforms previous methods for five different datasets under up to nine metrics. In future work, we will extend the selective auditory attention mechanism to the audio-only target speaker extraction. Furthermore, we will analyse the roles of each auxiliary reference to achieve an intelligent balance between them.